At 27 degree celsius, 0.8g of gas occupies 500 cm3 under normal pressure. What would be its volume? If the temperature is increased to 54 degree celsius at the same pressure.

Grade:11

1 Answers

Vikas TU
7 years ago
Hi,
Already answered your question plz check again:
At 27 degree celsius and 80 kilopascal pressure,
PV = nRT
0.789539 * 0.9 = 0.3*R*300/M
we getm
M = 90R/0.7105851 => 10.39847 g
At 127 ‘C
0.986923 * V = (0.3/10.39)*R*400
V = 0.960 Litre